OpenStudy (anonymous):

how does tan^4(x)=tan^2(x)sec^2(x)-tan^2(x)

OpenStudy (anonymous):

use the properties of trig functions sec2-1=tan^2 *right side* factor out tan^2 to get tan^2(sec^2-1)=tan^2tan^2 tan^4=tan^4

OpenStudy (anonymous):

should be sec^2-1=tan^2 not sec2-1=tan^2